# Risk-sensitive Reinforcement Learning Based on Convex Scoring Functions

### v1

@@ 6 Proofs
Case 2: $h(\cdot,\cdot)$ is constant valued in its second argument, i.e., $h(x,\upsilon)=\tilde{h}(x),$ and $f(y,\upsilon)=w(y)+\phi(\upsilon-\psi(y)),$ for some continuous functions $w,\phi,\psi.$ Let
@@
which leads to a contradiction with $\upsilon_{n}$ tending to infinity. Thus, we have $\bar{\upsilon}^{*}<\infty$ and $\underline{\upsilon}^{*}>-\infty.$ As $\upsilon\mapsto\mathbb{E}[f(Y,\upsilon)]$ is convex and $\upsilon^{*}$ is a minimizer of $\upsilon\mapsto\mathbb{E}[f(Y,\upsilon)]$ iff it is a minimizer of $\upsilon\mapsto\tilde{h}(\mathbb{E}[f(Y,\upsilon)]),$ we can define an interval $[m^{\underline{y},\bar{y}},M^{\underline{y},\bar{y}}]$ as follows: If $\bar{\upsilon}^{*},\underline{\upsilon}^{*}\in\mathbb{R},$ then we take $m^{\underline{y},\bar{y}}<\min\{\bar{\upsilon}^{*},\underline{\upsilon}^{*}\},$ $M^{\underline{y},\bar{y}}>\max\{\bar{\upsilon}^{*},\underline{\upsilon}^{*}\}$ and the conclusion follows; If $\bar{\upsilon}^{*}=-\infty$ and $\underline{\upsilon}^{*}=\infty,$ $\tilde{h}(\mathbb{E}[f(Y,\upsilon)])$ is constant of $\upsilon$ on $\mathbb{R}$ and any sub-interval of $\mathbb{R}$ satisfies; If $\bar{\upsilon}^{*}=-\infty$ and $\underline{\upsilon}^{*}\in\mathbb{R},$ then $\tilde{h}(\mathbb{E}[f(Y,\upsilon)])$ is constant of $\upsilon$ on $(-\infty,\underline{\upsilon}^{*}]$ and is increasing on $[\underline{\upsilon}^{*},\infty),$ and we take $m^{\underline{y},\bar{y}}<\underline{\upsilon}^{*},$ $M^{\underline{y},\bar{y}}>\underline{\upsilon}^{*}$; If $\bar{\upsilon}^{*}\in\mathbb{R}$ and $\underline{\upsilon}^{*}=\infty,$ then $\tilde{h}(\mathbb{E}[f(Y,\upsilon)])$ is constant of $\upsilon$ on $[\bar{\upsilon}^{*},\infty)$ and is increasing on $(-\infty,\bar{\upsilon}^{*}]$, and we take $m^{\underline{y},\bar{y}}<\bar{\upsilon}^{*},$ $M^{\underline{y},\bar{y}}>\bar{\upsilon}^{*}.$
@@
We define $\pi_{\upsilon,n}(\cdot|t,s,y)$ recursively as a point mass on $a_{n}.$ By the dominated control theorem, it is easy to see that
@@
$\left|\mathbb{E}[]f(Y,\upsilon_{1})]-\mathbb{E}[f(Y,\upsilon_{2})]\right|\leqslant\Delta|\upsilon_{1}-\upsilon_{2}|,$
@@
As $f(y,\upsilon)$ is convex in $\upsilon\in\mathbb{R},$ $L$ is increasing in $\upsilon_{2}$ and $\upsilon_{1}$ respectively, it is sufficient to show
@@
which then implies
@@
In what follows, we only show (20), and the proof for (19) is similar. Suppose (20) does not hold. Then there exist $\left\{y_{n}\right\}^{\infty}_{n=1}\subset\mathcal{Y}$ and $\upsilon_{n}\uparrow\bar{\upsilon}\ (n\to\infty)$ such that $f(y_{n},\bar{\upsilon})-f(y_{n},\upsilon_{n})>n(\bar{\upsilon}-\upsilon_{n}).$ Since $\underline{y}\leqslant y_{n}\leqslant\bar{y}$ for all $n\geqslant 1,$ there exists a convergent subsequence $y_{n_{k}}\to y_{0}\;(k\to\infty).$ Thus, we have
@@
First, we prove that $\upsilon\mapsto h(\mathbb{E}[V^{*}_{0,\upsilon}(S_{0},0)],\upsilon)$ is a continuous function on $\Upsilon.$ We shall prove this claim by contradiction, and suppose that $h(\mathbb{E}[V^{*}_{0,\upsilon}(S_{0},0)],\upsilon)$ has a discontinuous point at $\upsilon_{0}.$ Then, there exists a sequence $\left\{\upsilon_{n}\right\}_{n=1}^{\infty}$ satisfying $\upsilon_{n}\to\upsilon_{0}\ (n\to\infty)$ and
@@
for some $\delta>0.$ First, we note that
@@
If $\mathscr{U}^{*}=\Upsilon,$ the conclusion is trivial, so we only consider the case when $\mathscr{U}^{*}$ is a strict subset of $\Upsilon.$ We shall prove by contradiction and assume that there exist $\gamma_{0}>0$ and $\left\{(\phi_{n},\theta_{n})\right\}_{n=1}^{\infty}$ such that
@@
Because $\mathcal{S},\mathcal{Y},\Upsilon$ are compact and $f(\cdot,\cdot)$ is continuous, we have $|V_{0,\upsilon}(s_{0},0;\theta)|$ and $|V^{\phi}_{0,\upsilon}(s_{0},0;\theta)|$ are bounded for $(s_{0},\upsilon)\in\mathcal{S}\times\Upsilon$. Denote the upper bound by $B.$ If the third condition holds, we have
@@
The proof of Proposition 4 needs the following preparation.

### v2

@@ 6 Proofs
Case 2: $h(\cdot,\cdot)$ is constant with respect to its second argument, i.e., $h(x,\upsilon)=\tilde{h}(x),$ and $f(y,\upsilon)=w(y)+\phi(\upsilon-\psi(y)),$ for some continuous functions $w,\phi,\psi.$ Let
@@
which leads to a contradiction with $\upsilon_{n}$ tending to infinity. Thus, we have $\bar{\upsilon}^{*}<\infty$ and $\underline{\upsilon}^{*}>-\infty.$ As $\upsilon\mapsto\mathbb{E}[f(Y,\upsilon)]$ is convex and $\upsilon^{*}$ is a minimizer of $\upsilon\mapsto\mathbb{E}[f(Y,\upsilon)]$ iff it is a minimizer of $\upsilon\mapsto\tilde{h}(\mathbb{E}[f(Y,\upsilon)]),$ we can define an interval $[m^{\underline{y},\bar{y}},M^{\underline{y},\bar{y}}]$ as follows: If $\bar{\upsilon}^{*},\underline{\upsilon}^{*}\in\mathbb{R},$ then we take $m^{\underline{y},\bar{y}}<\min\{\bar{\upsilon}^{*},\underline{\upsilon}^{*}\},$ $M^{\underline{y},\bar{y}}>\max\{\bar{\upsilon}^{*},\underline{\upsilon}^{*}\}$ and the conclusion follows; If $\bar{\upsilon}^{*}=-\infty$ and $\underline{\upsilon}^{*}=\infty,$ then $\tilde{h}(\mathbb{E}[f(Y,\upsilon)])$ is constant of $\upsilon$ on $\mathbb{R}$ and the conclusion follows on any sub-interval of $\mathbb{R};$ If $\bar{\upsilon}^{*}=-\infty$ and $\underline{\upsilon}^{*}\in\mathbb{R},$ then $\tilde{h}(\mathbb{E}[f(Y,\upsilon)])$ is constant of $\upsilon$ on $(-\infty,\underline{\upsilon}^{*}]$ and is increasing on $[\underline{\upsilon}^{*},\infty),$ and we take $m^{\underline{y},\bar{y}}<\underline{\upsilon}^{*},$ $M^{\underline{y},\bar{y}}>\underline{\upsilon}^{*}$; If $\bar{\upsilon}^{*}\in\mathbb{R}$ and $\underline{\upsilon}^{*}=\infty,$ then $\tilde{h}(\mathbb{E}[f(Y,\upsilon)])$ is constant of $\upsilon$ on $[\bar{\upsilon}^{*},\infty)$ and is increasing on $(-\infty,\bar{\upsilon}^{*}]$, and we take $m^{\underline{y},\bar{y}}<\bar{\upsilon}^{*},$ $M^{\underline{y},\bar{y}}>\bar{\upsilon}^{*}.$
@@
We define $\pi_{\upsilon,n}(\cdot|t,s,y)$ recursively as a point mass on $a_{n}.$ By the dominated convergence theorem, it is easy to see that
@@
$\left|\mathbb{E}[f(Y,\upsilon_{1})]-\mathbb{E}[f(Y,\upsilon_{2})]\right|\leqslant\Delta|\upsilon_{1}-\upsilon_{2}|,$
@@
As $f(y,\upsilon)$ is convex in $\upsilon\in\mathbb{R},$ $L$ is increasing in $\upsilon_{2}$ and $\upsilon_{1}$ respectively. Hence, it is sufficient to show
@@
which then imply
@@
In what follows, we only show (20), and the proof for (19) is similar. We prove by contradiction and suppose that (20) does not hold. Then there exist $\left\{y_{n}\right\}^{\infty}_{n=1}\subset\mathcal{Y}$ and $\upsilon_{n}\uparrow\bar{\upsilon}\ (n\to\infty)$ such that $f(y_{n},\bar{\upsilon})-f(y_{n},\upsilon_{n})>n(\bar{\upsilon}-\upsilon_{n}).$ Since $\underline{y}\leqslant y_{n}\leqslant\bar{y}$ for all $n\geqslant 1,$ there exists a convergent subsequence $y_{n_{k}}\to y_{0}\;(k\to\infty).$ Thus, we have
@@
First, we prove that $\upsilon\mapsto h(\mathbb{E}[V^{*}_{0,\upsilon}(S_{0},0)],\upsilon)$ is a continuous function on $\Upsilon.$ We prove this claim by contradiction, and suppose that $\upsilon\mapsto h(\mathbb{E}[V^{*}_{0,\upsilon}(S_{0},0)],\upsilon)$ has a discontinuous point at $\upsilon_{0}.$ Then, there exists a sequence $\left\{\upsilon_{n}\right\}_{n=1}^{\infty}$ satisfying $\upsilon_{n}\to\upsilon_{0}\ (n\to\infty)$ and
@@
for some $\delta>0.$ We note that
@@
If $\mathscr{U}^{*}=\Upsilon,$ the conclusion is trivial. We only need to consider the case where $\mathscr{U}^{*}$ is a proper subset of $\Upsilon.$ We prove by contradiction and suppose that there exist $\gamma_{0}>0$ and $\left\{(\phi_{n},\theta_{n})\right\}_{n=1}^{\infty}$ such that
@@
Because $\mathcal{S},\mathcal{Y},\Upsilon$ are all compact and $f(\cdot,\cdot)$ is continuous, we have $|V_{0,\upsilon}(s_{0},0;\theta)|$ and $|V^{\phi}_{0,\upsilon}(s_{0},0;\theta)|$ are bounded for $(s_{0},\upsilon)\in\mathcal{S}\times\Upsilon$. Denote the upper bound by $B.$ If the third condition holds, we have
@@
The proof of Proposition 4 requires the following lemma.